### v1

# Generalizing fusion rules by shuffle: Symmetry-based classifications of nonlocal systems constructed from similarity transformations

Abstract
We study fusion rings, or symmetry topological field theories (SymTFTs), which lie outside the non-negative integer matrix representation (NIM-rep), by combining knowledge from generalized symmetry and that from pseudo-Hermitian systems. By applying the Galois shuffle operation to the SymTFTs, we reconstruct fusion rings that correspond to nonlocal CFTs constructed from the corresponding local nonunitary CFTs by applying the similarity transformations. The resultant SymTFTs are outside of NIM-rep, whereas they are ring isomorphic to the NIM-rep of the corresponding local nonunitary CFTs. We study the consequences of this correspondence between the nonlocal unitary model and local nonunitary models. We demonstrate the correspondence between their classifications of massive or massless renormalization group flows and the discrepancies between their boundary or domain wall phenomena. Our work reveals a new connection between ring isomorphism and similarity transformations, providing the fundamental implications of ring-theoretic ideas in the context of symmetry in physics.
pacs:
73.43.Lp, 71.10.Pm
YITP-25-181
I introduction
Similarity transformation is a fundamental transformations, which enables one to obtain the Hermitian systems from pseudo-Hermitian systems by the invertible transformations[1, 2, 3]. This also implements a dual relationship between the nonunitary conformal field theories (CFTs) and the nonunitary CFTs describing the corresponding pseudo-Hermitian or Hermitian lattice models respectively[4, 5]. One can see a similar (or the same) dual relationship in [6, 7, 8] related to the 3d3d3d-3d correspondence[9, 10, 11]. Related to the property of the effective vacuum or ground state, the Zamolodchikov $c$-theorem in unitary CFTs[12] has been generalized to $c_{\text{eff}}$-theorem[13] with a connection to the $PT$ symmetry in non-Hermitian systems[14, 15].
One can notice the nontrivial relationship between pseudo-Hermitian systems and the corresponding nonlocal Hermitian systems in the older studies on the realizations of chiral fermions. The lattice realization of the chiral fermions is restricted by the Nielsen-Ninomiya theorem[16, 17], and the two possible ways to realize chiral fermions are breaking locality or Hermiticity. There also exists a lattice analogue in $1+1$-dimensions[18]. We note a recent review on chiral fermion[19], those on non-Hermitian systems[20, 21, 22] as references.
We also note the pioneering nonlocal model, Haldane-Shastry model[23, 24] and several recent research directions on non-local systems, numerical studies [25, 26, 27], long-range interacting integrable models [28, 29], extension of Lieb-Schultz-Mattis theorem[30, 31, 32, 33, 34, 35], and quantum information scrambling[36, 37, 38, 39, 40, 41, 42, 43, 44]. These studies on non-local system has been developed in distinct research fields, with different techniques and languages. Hence, one may require some general method applicable to these models, such as symmetry analysis.
In this paper, we construct and study a particular type of conserved charges, symmetry topological field theories (SymTFTs)[45], of the nonlocal Hermitian systems derived from pseudo-Hermitian systems by the similarity transformations. For readers interested in SymTFTs and generalized symmetries, we note several earlier works[46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58], recent reviews[59, 60, 61, 62] and two works[63, 64] by the authors and a collaborator. The fundamental algebraic structures are a generalization of the Verlinde formula [65, 66, 67], Eq. (16), and the corresponding SymTFTs and their modules are written as Eq. (22) and Eq. (30). Their implications to RGs are summarized in Fig. 1. However, the defect or boundary phenomena are more subtle because of the nonlocality, as we demonstrate in Sec. V.
II Nonunitary CFTs and their symmetry
First, we concentrate our attention on the $A$-type diagonal CFTs with the following partition functions,
$Z=\sum_{\alpha}\chi_{\alpha}(\tau)\overline{\chi_{\alpha}}(\overline{\tau})$ (1)
where $\chi$ and $\overline{\chi}$ are the chiral or antichiral characters labelled by the primary field $\alpha$ and $\tau$ and $\overline{\tau}$ are the corresponding modular parameters. We note [68, 69, 70] as general references, and mainly follow the recent work[64] by the authors.
In a pseudo-Hermitian system, one can introduce the left and right energy eigenstates. Hence, corresponding to this fact, one can represent the characters as,
$\chi_{\alpha}(\tau)=\sum_{M}\widetilde{\langle\alpha,M|}e^{2\pi i\tau(L_{0}-c/12)}|\alpha,M\rangle$ (2)
where $\{|\alpha,M\rangle\}$ is the right eigenstates of the Virasoro generator $L_{0}(\in\{L_{m}\}_{m\in\mathbb{Z}})$ producing the corresponding Verma module labelled by the index $M$ of descendant fields, and the left eigenstates $\{\tilde{\langle\alpha|}\}$ is its linear dual satisfying the relation $\widetilde{\langle\alpha,m|}|\beta,n\rangle=\delta_{\alpha,\beta}\delta_{m,n}$. We also note the form of the Virasoro algebra
$L_{m},L_{n}]=(m-n)L_{m+n}+\frac{c}{12}(m^{3}-m)\delta_{m+n,0}$ (3)
where $c$ is the central charge and $\delta_{a,b}$ is the Kronekker delta. One can also implement the chiral projection operators $P_{\alpha,m}=|\alpha,m\rangle\widetilde{\langle\alpha,m|}$. In a similar way, one can also implement their antichiral analogue $\overline{P_{\alpha,\overline{M}}}$. Hence, we introduce the following idempotents,
$P_{\alpha}=\sum_{M,\overline{M}}P_{\alpha,M}\overline{P_{\alpha,\overline{M}}}$ (4)
satisyfying the relation $P_{\alpha}P_{\beta}=\delta_{\alpha,\beta}P_{\alpha}$. In this expression, the unitary CFTs can be interpreted as exceptional models where the linear dual coincides with the complex conjugation, i.e. $\left(|\alpha,M\rangle\right)^{\dagger}=\langle\alpha,M|$.
For the latter convenience, we introduce the modular $T$ and $S$ transformations,
$T:\tau\rightarrow\tau+1$ (5)
$S:\tau\rightarrow-1/\tau$ (6)
and their action to the characters,
$T:\chi_{\alpha}(\tau+1)=e^{2\pi i\left(h_{\alpha}-\frac{c}{24}\right)}\chi_{\alpha}(\tau)$ (7)
$S:\chi_{\alpha}\left(-1/\tau\right)=\sum_{\beta}S_{\alpha,\beta}\chi_{\beta}(\tau).$ (8)
The Verlinde formula[71] is written as follows:
$N^{\gamma}_{\alpha,\beta}=\sum_{\delta}\frac{S_{\alpha,\delta}S_{\beta,\delta}\overline{S_{\delta,\gamma}}}{S_{I,\delta}}$ (9)
where $I$ is the identity operator corresponding to the vacuum and $N$ is the so-called nonnegative integer matrix, specifying the fusion rule $\alpha\times\beta=\sum_{\gamma}N^{\gamma}_{\alpha,\beta}\gamma$. It should be stressed that even in nonunitary models, the matrix $N$ takes nonnegative integer values. In this setting, one can introduce the Verlinde line operator as follows[48],
$\mathcal{Q}_{\alpha}=\sum_{\beta}\frac{S_{\alpha,\beta}}{S_{I,\beta}}P_{\beta}$ (10)
where these operators also satisfy the relation $\mathcal{Q}_{\alpha}\times\mathcal{Q}_{\beta}=\sum_{\gamma}N^{\gamma}_{\alpha,\beta}\mathcal{Q}_{\gamma}$ by intepreting the symbol $\times$ as multiplication of operators. It should be remarked that $\mathbf{A}=\{\mathcal{Q}_{\alpha}\}_{\alpha}$ is as a set of linear operator over the complex number field $\mathbb{C}$ following the pioneering work[48]. By definition, these operators $\{\mathcal{Q}_{\alpha}\}$ defined by the projections commutes with the CFT Hamiltonian, and one can interprete them as conseved charges or SymTFTs. For latter use, we also note a relation,
$P_{\alpha}=\sum_{\beta}\mathcal{Q}_{\beta}S_{I,\alpha}\overline{S_{\alpha,\beta}}.$ (11)
Hence, $\{P_{\alpha}\}$ and $\{\mathcal{Q}_{\alpha}\}$ are ring isomorphic.
The form of $\mathcal{Q}_{\alpha}$ implies the existence of the modules (or smeared BCFTs[72, 73, 74, 75, 76, 63, 64]) spanned by the left and right Cardy’s states,
$|\alpha\rangle=\sum_{\beta}\frac{S_{\alpha,\beta}}{\sqrt{S_{I,\beta}}}|\beta\rangle\rangle,\widetilde{\langle\alpha|}=\sum_{\beta}\frac{\overline{S}_{\alpha,\beta}}{\sqrt{S_{I,\beta}}}\widetilde{\langle\langle\beta|}$ (12)
where $\{|\alpha\rangle\rangle\}$ is the Ishibashi states[77] satisfying the following relation,
$\widetilde{\langle\langle\alpha|}e^{\pi i\tau(L_{0}+\overline{L_{0}}-c/12)}|\beta\rangle\rangle=\chi_{\alpha}(\tau)\delta_{\alpha,\beta}.$ (13)
By applying the Verlinde formula, the symmetry operators act on the module in the following way[78],
$\mathcal{Q}_{\alpha}|\beta\rangle=\sum N_{\alpha,\beta}^{\gamma}|\gamma\rangle$ (14)
Moreover, the Cardy states satisfy the following relation[79, 80],
$Z_{\alpha,\beta}(\tau)=\widetilde{\langle\alpha|}e^{2\pi i(L_{0}+\overline{L_{0}}-c/12)}|\beta\rangle=\sum_{\gamma}N^{\gamma}_{\alpha,\beta}\chi_{\gamma}(-1/\tau)$ (15)
As can be seen in Eq.(9), the vacuum, $I$, plays the central role. However, in nonunitary CFTs, there exists another fundamental object, the effective vacuum, $o$. Hence, one can naively expect the following to have some meaning,
$N^{\gamma,[o]}_{\alpha,\beta}=\sum_{\delta}\frac{S_{\alpha,\delta}S_{\beta,\delta}\overline{S_{\delta,\gamma}}}{S_{o,\delta}}.$ (16)
This matrix is less familar, but it has already appeared in [65, 66], and the section $8.19.$ of [67], and it has been known to be integers. Moreover, we demonstrate that there exists the corresponding algebra, $\mathbf{A}^{[o]}=\{\mathcal{Q}_{\alpha}^{[o]}\}$, which is ring isomorphic to $\mathbf{A}=\{\mathcal{Q}_{\alpha}\}$. Hence, if there exists a unitary CFT which has the same character as a nonunitary CFT, one can obtain the dual relation between their symmetries. Fortunately, there exists at least two series of correspondence dual under the similarity, the correspondence between symplectic fermions and Dirac fermions[4] (or related $Sp(2n)_{k}\times Sp(2k)_{n}=SO(4nk)_{1}$ duality[81, 82]), and those between $M(2,2k+3)$ minimal models and $Osp(1,2)_{k}$ models (or related $W$-minimal models and $Osp(1,2n)_{k}$ models more generally[8]). We expect the same type of dualities for the models with positive $\{S_{o,\alpha}\}_{\alpha}$, and we write the corresponding SymTFTs and their modules.
III Characters and Symmetry operators by shuffle
First, we introduce the shuffle operation[83] $\eta^{[o]}$, the chiral analog of the similarity transformation[1, 2, 3], with effective-vacuum $o$ as follows,
$\eta^{[o]}:\eta^{[o]}(|\alpha,M\rangle)=|\alpha,M\rangle^{[o]}$ (17)
with the following shuffled representation of the Virasoro generator and the characters
$L_{0}^{[o]}-\frac{c_{\text{eff}}}{24}=\eta^{[o]}L_{0}{\eta^{[o]}}^{-1}-\frac{c}{24}$ (18)
$\chi_{\alpha}(\tau)=\langle\alpha,M|^{[o]}e^{2\pi i\tau(L_{0}^{[o]}-c^{[o]}/24)}|\alpha,M\rangle^{[o]}$ (19)
where $h_{\alpha,M}^{[o]}=h_{\alpha,M}-h_{o}$ and $c_{\text{eff}}=c^{[o]}=c-24h_{o}$ are called effective conformal dimensions or effective central charges, respectively. Significance of the above representations have been first noticed in [4] and applied to the analysis of gapless fractional quantum Hall states[84, 85, 86, 87] in [88, 89, 90, 76]. For the antichiral parts, we also assume the same shuffle operation and denote it as $\overline{\eta}^{[o]}$.
Whereas we assumed the existence of the shuffle operation, a chiral and antichiral analogue of similarity transformations, but the combination $\eta^{[o]}\overline{\eta}^{[o]}$ can rigorously be identified as the similarity tranformation by applying the arguments in [1, 2, 3]. Corresponding to the above unitary representation, we also introduce the following projections,
$P_{\alpha,M}^{[o]}=|\alpha,M\rangle^{[o]}\langle\alpha,M|^{[o]}$ (20)
We have assumed the Hermiticity for the quantum states, $\langle\alpha,M|^{[o]}=\left(|\alpha,M\rangle^{[o]}\right)^{\dagger}$, whereas they are defined by vector and linear dual in nonunitary CFTs. This is highly similar to the pseudo-Hermiticity [21]. By applying the same procedure for the antichiral parts, we introduce the following projection operators,
$P_{\alpha}^{[o]}=\sum_{M,\overline{M}}P_{\alpha,M}^{[o]}\overline{P_{\alpha,\overline{M}}^{[o]}}$ (21)
This expression induces the ring isomorphism $\{P_{\alpha}\}=\{P_{\alpha}^{[j]}\}$.
By definition, this shuffle operation induces the following symmetry operators,
$\mathcal{Q}_{\alpha}^{[o]}=\sum_{\beta}\frac{S_{\alpha,\beta}}{S_{o,\beta}}P_{\beta}^{[o]}$ (22)
with $S_{o,\beta}\neq 0$. In many series of nonunitary CFTs, the condition $\{S_{o,\beta}>0\}_{\beta}$ is satisfied. Hence, the above symmetry operators are well-defined in many cases, and one can obtain the expression,
$\mathcal{Q}_{\alpha}^{[o]}\times\mathcal{Q}_{\beta}^{[o]}=\sum_{\gamma}N^{\gamma,[o]}_{\alpha,\beta}\mathcal{Q}_{\gamma}^{[o]}$ (23)
by applying the definition of $N^{[o]}$, a generalization of Verlinde formula. Hence, we have demonstrated the implication of the fusion coefficient defined by the effective vacuum, only by assuming the unitarity. We also note that $\mathcal{Q}_{o}^{[o]}=\sum_{o}P_{\beta}^{[o]}$ behaves as the identity in $\mathbf{A}^{[o]}=\{\mathcal{Q}_{\alpha}^{[o]}\}$ and the roles of $I$ and $o$ are shuffled in this sense. In the next section, we demonstrate the naturalness of the algebra $\mathbf{A}^{[o]}$, by identifying their quantum dimensions and modules (or smeared BCFTs) and by studying their implications to RGs.
It should be stressed that the condition $\{S_{j,\beta}\neq 0\}_{\beta}$ implies the invertibility of $\mathcal{Q}_{j}$ in linear algebra. It is well-known that the invertibility of the symmetry plays a significant role in studying gauging operations[57, 58], and our work provides a generalization respecting the linear algebraic structure of quantum systems. Moreover, one can obtain the following relations,
$P_{\alpha}^{[o]}=\sum_{\beta}\mathcal{Q}_{\beta}^{[o]}S_{o,\alpha}\overline{S_{\alpha,\beta}}$ (24)
Hence, there exists a series of ring isomorphisms,
$\{\mathcal{Q}_{\alpha}\}=\{P_{\alpha}\}=\{P_{\alpha}^{[o]}\}=\{\mathcal{Q}_{\alpha}^{[o]}\}$ (25)
These relations imply the equivalence between the classifications of RGs (or SymTFTs) in nonunitary CFTs and those in nonlocal unitary CFTs. However, the matrix $N^{[o]}$ can be outside of NIM-rep[79, 80], and a simple object, such as $\mathcal{Q}_{\alpha}$, cannot be mapped to a simple object in $\mathbf{A}^{[o]}$ in general. For SymTFTs, this aspect does not affect the analysis, but it does the analysis for boundary conditions or defects, where the coefficients before objects need to become nonnegative integers. We also note that our generalization of fusion rules is different from those constructed from gauging operations[91, 92, 93, 94].
III.1 $M(2,5)$ minimal model and $Osp(1,2)_{1}$ WZW model
In the $M(2,5)$ minimal model, there exists only two primary fields $I$ and $o$ with the conformal dimensions $h_{I}=0$, $h_{o}=-1/5$, and with the central charge c=22/5c=-22/5 . The modular S matrix is given as follows,
$S=\frac{2}{\sqrt{5}}\begin{pmatrix}-\sin\left(\frac{2\pi}{5}\right)&\sin\left(\frac{4\pi}{5}\right)\\ \sin\left(\frac{4\pi}{5}\right)&\sin\left(\frac{2\pi}{5}\right)\end{pmatrix}$
The nontrivial part of the fusion ring $\mathbf{A}$ is,
$\mathcal{Q}_{o}\times\mathcal{Q}_{o}=\mathcal{Q}_{I}+\mathcal{Q}_{o},$ (26)
where $\mathcal{Q}_{I}$ is the identity operator. $\{\mathcal{Q}_{I},\mathcal{Q}_{o}\}$ is called the Fibonacci fusion ring and is famous for the description of a non-abelian anyon[95, 96].
By applying the shuffle operation, one can obtain the following fusion ring $\mathbf{A}^{[o]}$,
$\mathcal{Q}_{I}^{[o]}\times\mathcal{Q}_{I}^{[o]}=\mathcal{Q}_{o}^{[o]}-\mathcal{Q}_{I}^{[o]}$ (27)
where $\mathcal{Q}_{o}^{[o]}$ is the identity operator. As can be seen from the above expression, the roles of $I$ and $o$ are shuffled and their conformal dimensions are $h_{o}^{[o]}=0$, $h_{I}^{[o]}=1/5$. These conformal dimensions and the effective central charge $c_{\text{eff}}=2/5$ match with those of $Osp(1,2)_{1}$ Wess-Zumino-Witten (WZW) model[7, 8]. Hence, the above $\mathbf{A}^{[o]}$ provides a natural expression of SymTFT of the $Osp(1,2)_{1}$ WZW model.
IV Ring isomorphism and application to RGs
In this section, we study the consequences of the similarity transformation in SymTFTs combined with massive and massless RG. We note the ring isomorphism as $\iota:\mathbf{A}\rightarrow\mathbf{A}^{[o]}$ in the ultraviolet(UV) theories. At the algebraic level, the massless RG corresponds to a ring homomorphism (or functor), and this aspect has been studied in the name RG domain wall[97, 55, 98] or gapped domain wall[99, 100, 101, 102, 103]. On the other hand, the massive RG corresponds to the operation taking module (or smeared BCFT) of the unbroken symmetry subring of the original UV theory[104, 105, 106, 107, 72, 108, 73, 74, 75, 76, 109, 110, 63, 64].
IV.1 Massless RG flow
In this subsection, we note the implications of the similarity transformations realized as ring isomorphism. We denotes an infrared (IR) theories as $\mathbf{A^{\prime}}$ and $\mathbf{A^{\prime}}^{[o]}$ and their similarity transformation as $\iota^{\prime}:\mathbf{A}^{\prime}\rightarrow\mathbf{A^{\prime}}^{[o]}$.
First, we assume the massless RG or ring homomorphism between a UV theory $\mathbf{A}$ to an IR theory $\mathbf{A^{\prime}}$[100, 111, 103]. In this setting, one can implement the ring homomorphism of the corresponding nonlocal unitary model as
$\iota^{\prime}\rho\iota^{-1}:\mathbf{A}^{[o]}\rightarrow\mathbf{A^{\prime}}^{[o^{\prime}]}$ (28)
The unbroken symmetry $\mathbf{A}_{\text{ub}}$, such as $\mathbf{A}_{\text{ub}}\cap\text{Ker}\rho=\{0\}$ in the UV theory is transformed to $\iota(\mathbf{A}_{\text{ub}})$ and one can apply the analysis in [63, 64]. See FIG. 1 for intuition. For further studies, we introduce the following generalization of quantum dimensions (which can be called shuffled quantum dimensions),
$q_{\alpha,(a)}^{[o]}=\frac{S_{\alpha,a}}{S_{o,a}}.$ (29)
These quantum dimensions satisfy the relation $q_{\alpha,(a)}^{[o]}\times q_{\beta,(a)}^{[o]}=\sum_{\gamma}N^{\gamma,[o]}_{\alpha,\beta}q_{\gamma,(a)}^{[o]}$. Hence, one can implement the ring homomorphism $d_{(a)}^{[o]}:\mathbf{A}\rightarrow\mathbb{C}$ as $d_{(a)}^{[o]}(\alpha)=S_{\alpha,a}/S_{o,a}$.
Interestingly, the simple objects in $\mathbf{A}_{\text{ub}}$ are not mapped to simple objects in $\iota(\mathbf{A}_{\text{ub}})$ in general. By combining this observation with the homomorphism $\rho$, one can notice that the nonsimple (or nonlocal) objects in $\mathbf{A}$ can be simple objects in the IR theory $\mathbf{A^{\prime}}_{\text{ub}}^{[o]}$. At the level of linear algebra or ring theory, this phenomenon is understandable, but its phenomenologies or categorical descriptions are less clear.
[…]
Figure 1: The schematic correlation chart of the massless RG flow from UV theory to IR theory. We use "" symbol for IR rings. Notice that the RG flow is generally complicated in the sense that the simple objects in UV can be transformed to some complicated objects in IR.
IV.2 Smeared BCFT and massive RG flow
Corresponding to the representation of the symmetry operators, one can implement the following as smeared Cardy states $\{|\alpha\rangle^{[o]}\}$ and their complex conjugations $\{\langle\alpha|^{[o]}\}$,
$|\alpha\rangle^{[o]}=\sum_{\beta}\frac{S_{\alpha,\beta}}{\sqrt{S_{o,\beta}}}|\beta\rangle\rangle^{[o]},\langle\alpha|^{[o]}=\sum_{\beta}\frac{\overline{S}_{\alpha,\beta}}{\sqrt{S_{o,\beta}}}\langle\langle\beta|^{[o]}$ (30)
where the Ishibashi states $\{|\alpha\rangle\rangle\}$ satisfy the relation $\langle\langle\alpha|^{[o]}e^{\pi i\tau H}|\beta\rangle\rangle^{[o]}=\chi_{\alpha}(\tau)\delta_{\alpha,\beta}$. We stress the fundamental significance of the condtion $\{S_{o,\alpha}>0\}_{\alpha}$ in constructing the bra and ket by the complex conjugation, in comparison with nonunitary CFTs.
These boundary states can be interpreted as module of $\mathbf{A}^{[o]}=\{\mathcal{Q}_{\alpha}^{[o]}\}$, satisfying the following relation[78],
$\mathcal{Q}_{\alpha}^{[o]}|\beta\rangle^{[o]}=\sum_{\gamma}N^{\gamma,[o]}_{\alpha,\beta}|\gamma\rangle^{[o]}\ \left(=|\alpha\times\beta\rangle^{[o]}\right).$ (31)
Hence, by assuming the massive RG is realized as reduction of the symmetry from $\mathbf{A}^{[o]}$ to $\mathbf{A}_{\text{ub}}^{[o]}$[72, 73, 74, 75, 76, 63, 64], one can classify the gapped phase as $\mathbf{A}_{\text{ub}}^{[o]}$ invariant subspace of the Hilbert space spaned by $\{|\alpha\rangle^{[o]}\}$.
In this section, the unitarity combined with recent developments in BCFT[78, 72] has played a fundamental role, and the existing arguments in the smeared BCFT is applicable to our cases. However, the Cardy’s condition which determines the boundary critical phenomena is broken, because the amplitude can become outside of NIM-rep[80],
$\begin{split}Z_{\alpha,\beta}^{[o]}(\tau)&=\langle\alpha|^{[o]}e^{\pi i\tau(L_{0}+\overline{L_{0}}-c_{\text{eff}}/12)}|\beta\rangle^{[o]}\\ &=\sum_{\gamma}N^{\gamma,[o]}_{\alpha,\beta}\chi_{\gamma}(-1/\tau)\end{split}$ (32)
[…]
V Difficulty in boundary or defect phenomena
As we have observed in the previous section, the ring $\mathbf{A}^{[o]}$ outside of the NIM-rep[79, 80] cannot describe the defect or boundary phenomena straightforwardly. However, there still exist similarity transformations for the non-Hermitian system with such objects. For simplicity, we concentrate our attention on the annulus nonunitary CFT determined by $Z_{\alpha,\beta}$.
In this setting, by the open-closed duality[80], one can map the Hilbert space to that of the chiral CFT, $\oplus_{\gamma}N^{\gamma}_{\alpha,\beta}\gamma$. Hence, we propose a conjecture:
[…]
Under this assumption, one can apply the existing discussions on the corresponding vertex-operator algebraic arguments to the unexplored boundary physics of the nonlocal models. More technically, the doubling trick[113, 114] will play a fundamental role in studying the correlation functions and related quantities. We also point out that the above similarity transformation induced by some invertible operator can be regarded as a generalization of the large gauge transformation induced by unitary operators, known as Lieb-Schultz-Mattis operators[115, 116, 117, 118] [^1] We thank Masaki Oshikawa for the corresponding discussion.. This interpretation has appeared in [120].
By choosing the left and right boundary conditions carefully for $Z_{\alpha,\beta}^{[o]}$, one can also avoid the negative amplitude. This nontrivial (or nonlocal) relation between the nonnegative amplitude condition and the boundary conditions can be interpreted as a signal of the nonlocality. However, to our knowledge, the relationship between this nonlocality and the similarity transformation has not been studied. For example, in $M(2,5)$ minimal model, one can obtain the same characters from the pairs $(Z_{o,o}^{[o]},Z_{I,o})$, $(Z_{o,I}^{[o]},Z_{I,I})$, but it is impossible to reproduce $Z_{o,o}$ from $Z^{[o]}_{\alpha,\beta}$.
The phenomenon seems to be related to the boundary sensitity[121, 122, 123] or quantum skin effect[124] in non-Hermitian systems, but their precise relationship is still unclear. We also note some works studying similar problems arising from the boundary conditions of pseudo-Hermitian systems[125, 126], and a pioneering work pointing out its relation to quantum groups[127].
VI Conclusion and discussion
In this paper, we have studied the implication of the Galois shuffle operations in CFTs[83] and their implications as similarity transformations between pseudo-Hermitian systems and the corresponding nonlocal Hermitian systems. Essentially, this is a ring isomorphism between two theories, and the RGs of one side are compatible with the other side. However, when studying the implications for defects or boundaries, more careful treatment is necessary, and this difficulty itself is a natural consequence of the nonlocality.
[…]
VII Acknowledgement
The authors thank Naomichi Hatano, Hosho Katsura, and Po-Yao Chang for sharing the knowledge on the similarity transformations. YF thanks Yuma Furuta and Shinichiro Yahagi for related collaborations. YF also thanks Ingo Runkel and Sylvain Ribault for the discussions on fusion rules and Jurgen Fuch and Yuji Tachikawa for the helpful discussions on fusion ring. We thank Masaki Oshikawa for the helpful discussions on the large gauge transformation. TK is supported by Grant-in-Aid for JSPS Fellows No. 23KJ1315. YF thanks support from NCTS.
Appendix A Generalization and conjecture on similarity transformation
In this section, we comment on a straightforward generalization of our discussion by replacing the symbol $[o]$ to the general symbol $[j]$, satisfying the relation $S_{j,\alpha}\neq 0$ for all $\alpha$. This condition itself is nothing but the invertibility condition of the operator $\mathcal{Q}_{j}$ or anyon $j$ in the fusion ring. Hence, we conjecture
$\text{Invertible anyon $j$}\sim\text{similarity transformation $\eta^{[j]}$}$ (35)
However, we note that the $\mathcal{Q}_{j}$ itself does not implement the similarity transformation because it commute with Hamiltonian. Because one can apply the method in the main text only by replacing the symbols (and complex conjugate to linear dual), we only note the fusion coefficients:
$N^{\gamma,[j]}_{\alpha,\beta}=\sum_{\delta}\frac{S_{\alpha,\delta}S_{\beta,\delta}\overline{S_{\delta,\gamma}}}{S_{j,\delta}}.$ (36)
In Hermitian systems, it is known that the Lieb-Shultz-Mattis operators $U_{\alpha}$ which can be realized as an unitary operation changes or shuffle the energy eigenstates. Moreover, they correspond to primary field $\alpha$ and generate many-body excitation from the ground state $U_{\alpha}|0\rangle=\alpha(0)|0\rangle=|\alpha\rangle$ where we have assumed the radial quantization in CFTs and $|0\rangle$ is the vacuum. This unitary operation is called a large gauge transformation and the analogy between the similarity transformation and the gauge transformation has already been pointed out in[120]. Hence, we conjecture the similarity transformation can be implemented as invertible operator $U_{j}$ corresponding to the invertible anyon $j$ or the many-body excitation via the CFT expression $U_{j}|0\rangle=j(0)|0\rangle=|j\rangle$ where $j(0)$ corresponds to a generalized version of LSM operator $U_{j}$ generating the many-body excitations[115, 116, 117, 118].